OpenStudy (anonymous):

Convert the Cartesian equation x^4 - y^4 = 16 to a polar equation? r^2 = 16/cos2θ r^4 = 16/cos2θ r^2 = 8

OpenStudy (anonymous):

so far i have r^4(cos^4(theta)-sin^4(theta))=16

OpenStudy (anonymous):

Do you know any trig identities to simplify that?

OpenStudy (anonymous):

I think the answer is B, but i'm not totally sure

OpenStudy (anonymous):

I don't know I just learned this today and my worksheet doesn't say anything about it.

OpenStudy (anonymous):

But I think your right.

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

I think it simplifies to cos(2x)(1)

OpenStudy (anonymous):

r^4cos(2x)=16 r^4=16/cos(2x)

OpenStudy (anonymous):

Ok thanks!

OpenStudy (anonymous):

https://www.youtube.com/watch?v=JSBUnq535vM

OpenStudy (anonymous):

that explains the identity

OpenStudy (anonymous):

no problem